OpenStudy (anonymous):

The lateral area of a cone is 500 pi cm.^2. The radius is 50 cm. Find the slant height to the nearest tenth a. 10 cm b. 9.1 cm c. 14.2 cm d. 11.2 cm

OpenStudy (anonymous):

lateral are of a cone is LA = pi * r * s 500 Pi = Pi * 50 * s 500 = 50 * s 10 = s

OpenStudy (anonymous):

Thanks!